# Don’t You (Project Around Discs)? Neural Network Surrogate and Projected Gradient Descent for Calibrating an Intervertebral Disc Finite Element Model

### v1

@@ 3.1.1 Training set size and model architecture
Elapsed simulation time (hours) 29.6 118.4 236.9
As shown in Table 1, increasing the dataset size significantly improves most models. For instance, the RF model’s MAE decreases from 0.65 to 0.40 as the dataset grows from 128 to 1024, demonstrating that larger datasets help models capture underlying patterns more effectively. Simpler models like LR and SVR struggle with the non-linear interactions in the FE simulation [7], resulting in poor performance with MAEs of 1.05 and 1.04, respectively, even when trained on 1024 samples. This makes them less suitable for accurate predictions in this context. The NN model considerably improves, achieving an MAE of 0.10 and $\overline{\mathcal{R}^{2}}$ of 0.99 for 1024 samples. Notably, the NN model’s MAE of 0.43 for 128 samples may already be sufficient for some applications, requiring only about 30 hours to create the dataset. Observing that the NN surrogate begins to converge at 512 samples, we selected the best-performing NN model, trained on 1024 samples, for all subsequent experiments.
@@ 3.1.2 Interpolation and extrapolation
The ability to interpolate and extrapolate to unseen moments is crucial for the developed calibration method, as it relies on surrogate models and experimental measurements that are often reported at moments different from those used for training the surrogate. [^3] For instance, Heuer et al. [9] published their measurements for moments (1.25, 2.5, 3.75, 5, 6.25, 7.5, 10) Nm. These generalization abilities are evaluated in Fig. 4, where the top four performing surrogate models are tested on a dataset of 64 synthetic configurations with ground-truth RoM values for moments in the range (0.5, 10) Nm at 0.5 intervals. The evaluation assesses performance on moments on which the model was trained (1, 2, 3, 4, 5), interpolated moments (1.5, 2.5, 3.5, 4.5), and extrapolated moments (0.5 and $>5$). Cross-validated results are provided in Table 2.
@@
Table 2: Comparison of interpolation and extrapolation ability of the top performing surrogate models. The row colors match those of Fig. 4: white for trained moments, peach for moment extrapolation, and blue for moment interpolation. The results are the mean and standard deviation across the folds.
@@ 3.2 FE model calibration
In this section, the performance of the proposed calibration method is compared to the baselines. First, we test the calibration methods on a synthetic dataset to establish baseline performance and computational efficiency. We apply the calibration techniques to in vitro experimental measurements from Nicolini et al. [44] and Heuer et al. [9], assessing their practical applicability and robustness.
@@ 3.2.1 Synthetic data
Total running time (seconds) 2.58 11.92 0.01
Table 3 demonstrates that PGD w/ NN outperforms both baselines on synthetic data, achieving the lowest MAE of 0.06 and an $\overline{\mathcal{R}^{2}}$ score of 0.99, with a runtime of approximately 2.6 seconds. The GA w/ NN method also performs well, with an MAE of 0.18 and $\overline{\mathcal{R}^{2}}$ of 0.96, but it exhibits a more considerable variability across moments for most load cases. The high performance of both methods is expected since the calibrated target RoM values were derived from the same FE model used to create the training dataset. Though the inverse model is the fastest, requiring only 0.01 seconds, it fails to obtain applicable configurations, yielding a mean MAE of 0.52 and $\overline{\mathcal{R}^{2}}$ of 0.54. The primary issue is the ill-posedness of the inverse setting [46], where numerous, very different configurations can result in almost identical RoM measurements. Due to this lack of convergence, we did not evaluate the inverse model on experimental data.
@@ 3.2.2 Experimental specimen
To asses their relevance to real-world problems, we evaluated the calibration methods on six in vitro experimental specimens. We used five specimens reported by Nicolini et al. [44] in the same conditions used to train the surrogate. Additionally, a mean of eight specimen measurements reported by Heuer et al. [9] was utilized and preprocessed as suggested by Gruber et al. [14]. This measurement is available at moments different from the ones for which the surrogate was trained. Each method was given the target RoM values and tasked with predicting the corresponding configurations. Different from before, these predicted configurations were then validated with FE simulations to assess the agreement between the simulated and target RoM.
Figure 5: Example calibration results for samples 1 (top) and 3 (bottom) from Nicolini et al. [44]. The plots show the RoM values for the in vitro measurements (blue), the calibrated configuration using the proposed method (orange), and the GA w/ NN baseline (green). The proposed method performs better on sample 1 than sample 3. The poorer performance on sample 3 is attributed to the large variability in RoM across load cases in the target configuration, which differs significantly from the surrogate’s training data. Results for the other specimens are in C.
Detailed examples are shown in Fig. 5 and Table 4 summarizes the results. The GA baseline required up to 47 generations to converge on the specimens (see C). Notably, running the GA with FE simulations instead of the NN surrogate on the same machine would have taken up to 8 days per specimen for comparable results to our method, which is impractical for this study’s time constraints.
@@
Fig. 6 analyzes the error values across moments, averaged across load cases for experimental data. A significant increase in error is apparent at the lowest moment for many specimens. However, the surrogate evaluation in Sec. 3.1.2 suggests this error does not originate from the surrogate itself. An increase in error is also noticeable at the highest moment. We hypothesize that many measurements, such as the first sample by Nicolini et al. [44], exhibit high non-linearity, particularly in the transition between 1 and 2 Nm. This high non-linearity challenges the underlying FE model, producing poorer calibration results. Observations about the difficulties of this FE model in capturing diverse specimens while assuming identical geometry were reported by Nicolini et al. [44, 7].
Figure 6: Comparison of MAE vs. moment averaged across load cases on experimental specimens. Specimens are color-coded. The proposed method outperforms the baseline except for the third sample from Nicolini et al. [44] (purple). Both methods present an increased error at the lowest and highest moments.
@@
We hypothesize that this substantial variability in experimental data negatively impacts the optimization process. The GA w/ NN baseline, better suited for global optimization [47], excels in exploring a wider search space and finding configurations outside the training distribution. However, it still required five times more generations to converge for the third sample from Nicolini et al. [44] compared to the other specimens. [^4] The GA w/ NN required 47 generations to converge to an $\overline{\mathcal{R}^{2}}$ higher than 0.9 for this specimen vs 9.2 generations on average for the rest. See C. In contrast, our method relies on NN weights derived from the training set, which do not account for such extreme configurations.
3.2.3 Ablation studies
@@
The results demonstrate that the projection step is essential for ensuring feasible calibration outcomes. In the “Adam, L1 loss” setting, while the $\overline{\mathcal{R}^{2}}$ is highest at 0.95 and the MAE is lowest at 0.04, it results in a sum of 118.38 parameter values exceeding the bounds, making them infeasible for FE model input. With projection applied (“Adam, L1 loss, projection (Ours)”), the exceeding values drop to zero, though with a slight decrease in performance metrics ($\overline{\mathcal{R}^{2}}$ of 0.85 and MAE of 0.21).
@@
The “Adam, L1+bounds loss” setting, incorporating a penalty for out-of-bound values similar to Maso Talou et al. [33], reduces infeasible values to 0.11. However, it does not eliminate them, rendering some results unusable for the FE model. Not listed here, higher penalty weights reduced exceeding values but significantly degraded performance. Using an L2 loss (“Adam, L2 loss, projection”) treats large errors more severely, encouraging the minimization of significant deviations. An NN surrogate trained with L2 loss was used for consistency in this setting. However, this loss performed slightly worse than L1 loss, with an $\overline{\mathcal{R}^{2}}$ of 0.84 and an MAE of 0.25. Lastly, following Maso Talou et al. [33], we tried the L-BFGS optimizer instead of Adam. The “L-BFGS, L1 loss, projection” setting performed poorly, with an $\overline{\mathcal{R}^{2}}$ of -1.0 and an MAE of 0.93, indicating that this optimizer is unsuitable for this task.
@@ 4 Conclusions and outlook
This study introduced a novel calibration method for an FE model of the human L4-L5 IVD guided by an NN surrogate. Using an NN surrogate provides near-instant predictions, allowing for rapid calibration through optimization with PGD. Additionally, incorporating the projection step ensures optimized parameter values remain within feasible bounds, addressing a potential pitfall in existing NN surrogate optimization-based calibration methods. This balance between performance and feasibility is crucial for practical applications in clinical settings. Moreover, the developed method outperformed the evaluated baselines. While the GA w/NN baseline was effective, it required more time and relied on an external algorithm. Despite its rapid inference, the inverse model struggled with ill-posedness, failing to produce useful material configurations.
However, the study revealed limitations in our approach and the underlying FE model. The high variability in moment magnitude among certain experimental specimens led to sub-optimal convergence, as the NN surrogate was not trained on such extremes. Augmenting the training dataset with more diverse cases could enhance robustness. Future work should focus on generating representative samples using active learning [48], which involves selectively querying the FE model to include challenging, real-world scenarios. Furthermore, incorporating varying geometric parameters as FE model inputs, rather than fixed average dimensions, could allow surrogate models to be used for patient-specific simulations, resulting in more accurate calibration and considerable applicability. Another potential extension is adapting the surrogate model for a multi-body environment, as suggested by Hammer et al. [17].
@@ Acknowledgments
Funding: This work was supported by the the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program. Grant no: 101045128-iBack-epic-ERC-2021-COG
@@ B.2 Calibration baselines
The GA [7, 14] was configured with a population size of 20, evolving from initially sampled random configurations within the parameter ranges. The six best-performing individuals from each generation are selected for Crossover and Mutation. During Crossover, four new individuals are created by combining pairs of the selected individuals. Mutation generates four new individuals by randomly altering a parameter of one of the selected individuals. Additionally, Immigration is performed, introducing six new individuals per generation. The process iterates until an $\overline{\mathcal{R}^{2}}=1$ stopping criteria is met or $100$ generations are reached.
@@ Appendix C Detailed results
Figure 8: GA evolution of $\overline{\mathcal{R}^{2}}$ values for experimental specimens. 3rd sample by Nicolini et al. [44] requires more than double the amount of generations to converge to a score higher than 0.9.
@@
Appendix D List of abbreviations

### v2

@@ 3.1.1 Model architecture
Polynomial Regression (PR), degree 2 0.62$\pm$0.07 0.96$\pm$0.23 0.68$\pm$0.08 0.85$\pm$0.21 0.70$\pm$0.09 0.81$\pm$0.22
degree 3 0.62$\pm$0.05 0.94$\pm$0.19 0.79$\pm$0.07 0.68$\pm$0.19 0.82$\pm$0.07 0.61$\pm$0.22
degree 4 0.53$\pm$0.16 0.99$\pm$0.11 0.77$\pm$0.09 0.70$\pm$0.10 0.88$\pm$0.03 0.49$\pm$0.11
degree 5 0.45$\pm$0.21 1.07$\pm$0.11 0.74$\pm$0.10 0.75$\pm$0.10 0.86$\pm$0.03 0.53$\pm$0.10
@@
Dataset creation time (days) 1.2 4.9 9.8
3.1.2 Training set size
The size of the dataset used for training surrogate models is critical due to the time-consuming nature of FEM simulations. Table 1 explores the number of samples required for accurate model predictions, comparing datasets of 128, 512, and 1024 samples. Increasing the dataset size significantly improves most models. For instance, the RF model’s MAE decreases from 0.65 to 0.40 as the dataset grows from 128 to 1024, demonstrating that larger datasets help models capture underlying patterns more effectively. Notably, the NN model’s MAE of 0.43 for 128 samples may already be sufficient for some applications, requiring only about 1.2 days to create the dataset. Observing that the NN surrogate begins to converge at 512 samples, we selected the best-performing NN model, trained on 1024 samples.
@@ 3.1.3 Interpolation and extrapolation
The ability to interpolate and extrapolate to unseen moments is crucial for the developed calibration method, as it relies on surrogate models and experimental measurements that are often reported at moments different from those used for training the surrogate. [^3] For instance, Heuer et al. [15] published their measurements for moments (1.25, 2.5, 3.75, 5, 6.25, 7.5, 10) Nm. These generalization abilities are evaluated in Fig. 4, where the top four performing surrogate models are tested on a dataset of 64 synthetic configurations with ground-truth RoM values for moments in the range (0.5, 10) Nm at 0.5 intervals. The evaluation assesses performance on moments on which the model was trained (1, 2, 3, 4, 5), interpolated moments (1.5, 2.5, 3.5, 4.5), and extrapolated moments (0.5 and $>5$). Cross-validated results are provided in C.
@@ 3.2 FE model calibration
In this section, the performance of the proposed calibration method was compared to the baselines. First, we tested the calibration methods on a synthetic dataset to establish baseline performance and computational efficiency. We then applied the calibration techniques to in vitro experimental measurements conducted by Nicolini et al. [39] and Heuer et al. [15], assessing their practical applicability and robustness.
@@ 3.2.1 Synthetic data
Calibration time (seconds) 2.58 11.92 0.01
Table 2 demonstrates that PGD w/ NN outperforms both baselines on synthetic data, achieving the lowest MAE of 0.06 and an $\overline{\mathcal{R}^{2}}$ score of 0.99, with a runtime of approximately 2.6 seconds. The GA w/ NN method also performs well, with an MAE of 0.18 and $\overline{\mathcal{R}^{2}}$ of 0.96, but it exhibits a more considerable variability across moments for most load cases. The high performance of both methods is expected since the calibrated target RoM values were derived from the same FE model used to create the training dataset. Though the inverse model is the fastest, requiring only 0.01 seconds, it fails to obtain applicable configurations, yielding a mean MAE of 0.52 and $\overline{\mathcal{R}^{2}}$ of 0.54.
The inverse setting requires mapping RoM measurements back to distinct material input configurations, which is highly challenging. Many different configurations can produce nearly identical RoM values due to the non-monotonicity of the response surface [46], resulting in the ill-posedness of the inverse problem [58]. Given this lack of convergence, we did not proceed with evaluating the inverse model on experimental data.
@@ 3.2.2 Experimental specimen
To asses their relevance to real-world problems, we evaluated the calibration methods on six in vitro experimental specimens. We used five specimens from experiments conducted by Nicolini et al. [39], reported in the same conditions used to train the surrogate. Additionally, a mean of eight specimen measurements reported by Heuer et al. [15] was utilized and preprocessed as suggested by Gruber et al. [20]. This measurement is available at moments different from the ones for which the surrogate was trained. Each method was given the target RoM values and tasked with predicting the corresponding configurations. Different from before, these predicted configurations were then validated with FE simulations to assess the agreement between the simulated and target RoM. Example results are plotted in Fig. 5, Table 3 provides a summary, and the calibrated configurations are given in C.
Figure 5: Example calibration results for samples 1 (top) and 3 (bottom) from experiments by Nicolini et al. [39]. The plots show the RoM values for the in vitro measurements (blue), the calibrated configuration using the proposed method (orange), and the GA w/ NN baseline (green). The proposed method performs better on sample 1 than on sample 3. The poorer performance on sample 3 is attributed to the large variability in RoM across load cases in the target configuration, which differs significantly from the surrogate’s training data. See C for the other specimens.
@@
Fig. 6 analyzes the error values across moments, averaged across load cases for experimental data. A significant increase in error is apparent at the lowest moment for many specimens. However, the surrogate evaluation in Sec. 3.1.3 suggests this error does not originate from the surrogate itself. An increase in error is also noticeable at the highest moment. We hypothesize that many measurements, such as the first sample by Nicolini et al. [39], exhibit high non-linearity, particularly in the transition between 1 and 2 Nm. This high non-linearity challenges the underlying FE model, producing poorer calibration results. Observations about the difficulties of the FE model in capturing diverse specimens while assuming identical geometry were reported by Nicolini et al. [39, 10].
Figure 6: Comparison of MAE vs. moment averaged across load cases on calibrated experimental specimens. Specimens are color-coded. The proposed method outperforms the baseline except for the third sample from Nicolini et al. [39] (purple). Both methods present an increased error at the lowest and highest moments.
@@
We hypothesize that this substantial variability in experimental data negatively impacts the optimization process. The GA w/ NN baseline, better suited for global optimization [59], excels in exploring a wider search space and finding configurations outside the training distribution. However, it still required five times more generations to converge for the third sample from Nicolini et al. [39] compared to the other specimens. [^4] The GA w/ NN required 47 generations to converge to an $\overline{\mathcal{R}^{2}}$ higher than 0.9 for this specimen vs 9.2 generations on average for the rest. In contrast, our method relies on NN weights derived from the training set, which do not account for such extreme configurations.
Computational efficiency
We compared the efficiency of the proposed method with estimated run-times for a traditional GA that directly uses FE simulations rather than the surrogate. [^5] Details of how we obtained these estimates are provided in D. In Table 4, our method requires dataset generation and model training, but these are performed once, and calibration is parallelized, making its time independent of the number of specimens $n$. In contrast, while the GA does not require preliminary steps, its overall running time increases linearly with $n$. PGD w/ NN becomes more efficient than the GA already when calibrating more than one specimen in the worst case or more than three specimens in the average case.
Method Preliminary steps Calibration End-to-end
once per-specimen $n$ specimens
PGD w/ NN (Ours) 9.8 days (dataset + training) 2.58 sec 9.8 days
Traditional GA (Estimate) Avg. case N/A 2.7 days $2.7\times n$ days
Worst case 7.9 days $7.9\times n$ days
Table 4: Running time comparison between the developed method and average and worst-case estimates for a traditional GA. The running time of the developed method is independent of the number of calibrated specimens.
3.3 Validation
To validate the calibrated configurations we compared the simulated IDP with reference ranges reported by Brinckmann and Grootenboer [55], who provided experimental IDP measurements under axial compression for L4-L5 IVD specimens under three applied compressive forces: 300, 1000, and 2000 N. The analysis in Fig. 8 shows that all calibrated configurations fall within the reference ranges, demonstrating that the model accurately replicates the in vitro values observed in several human IVD specimens, ensuring the reliability of the calibration.
Figure 8: Model validation of calibrated specimens with IDP under axial compression. The black horizontal bars indicate the mean, minimum, and maximum measurements in the literature [55]. The colored lines show the simulation results for the calibrated specimens configurations, all within the reference ranges.
3.4 Ablation studies
@@
The results demonstrate that the projection step is essential for ensuring feasible calibration outcomes. In the “Adam, L1 loss” setting, while the $\overline{\mathcal{R}^{2}}$ is highest at 0.95 and the MAE is lowest at 0.04, it results in a sum of 118.38 parameter values exceeding the bounds, making them infeasible for FE model input. With projection applied (“Adam, L1 loss, projection (Ours)”), the exceeding values drop to zero, though with a decrease in performance metrics ($\overline{\mathcal{R}^{2}}$ of 0.85 and MAE of 0.21).
@@
The “Adam, L1+bounds loss” setting, incorporating a penalty for out-of-bound values similar to Maso Talou et al. [38], reduces infeasible values to 0.11. However, it does not eliminate them, rendering some results unusable for the FE model. Not listed here, higher penalty weights reduced exceeding values but significantly degraded performance. Using an L2 loss (“Adam, L2 loss, projection”) treats large errors more severely, encouraging the minimization of significant deviations. An NN surrogate trained with L2 loss was used for consistency in this setting. However, this loss performed slightly worse than L1 loss, with an $\overline{\mathcal{R}^{2}}$ of 0.84 and an MAE of 0.25. Lastly, we tried the L-BFGS optimizer instead of Adam. The “L-BFGS, L1 loss, projection” setting performed poorly, with an $\overline{\mathcal{R}^{2}}$ of -1.0 and an MAE of 0.93, indicating that this optimizer is unsuitable for this task. While these experiments highlight the strength of PGD in our setting, alternative methods for constrained optimization could be explored in future studies, such as Lagrange multipliers [60] or recent enhancements of PGD [61].
@@ 4 Conclusions and outlook
This study introduced a novel calibration method guided by an NN surrogate. Using an NN surrogate provides near-instant predictions, allowing for rapid calibration through optimization with PGD. Additionally, incorporating the projection step ensures optimized parameter values remain within feasible bounds, addressing a potential pitfall in existing NN surrogate optimization-based calibration methods. This balance between performance and feasibility is crucial for practical applications in clinical settings. Moreover, the developed method outperformed the evaluated baselines. While the GA w/NN baseline was effective, it required more time and relied on an external algorithm. Despite its rapid inference, the inverse model struggled with ill-posedness, failing to produce useful material configurations. We conducted our study on an FE model of the human L4-L5 IVD, future work could extend to other biomechanical domains, as we believe our method can be adapted to other FE models that face similar challenges in calibration.
Surrogate modeling
High variability in moments caused sub-optimal convergence, as the NN was not trained on such extremes. Augmenting the dataset with diverse cases could improve robustness. Future work should explore active learning [62] to generate representative samples by selectively querying the FE model for challenging, real-world scenarios. While the NN’s superior accuracy is beneficial, its limited explainability remains a concern in clinical applications. Future studies should explore methods to understand its internal decision-making process [63], such as the feature importance approach we used in Gruber et al. [46]. Our findings there could also help reduce the surrogate’s complexity by removing less impactful inputs. Finally, the GA baseline was evaluated with the NN surrogate rather than FE simulations. Despite precautions like setting a high GA stopping criterion and the NN’s high accuracy, some approximation error is expected.
Patient-specific clinical applications
Patient-specific models must consider disc material properties and geometric variations, especially in diseased discs with significant structural changes [64, 65]. Medical imaging, e.g., MRI, provides a promising path to integrate specific geometry into future models [66, 67, 68]. Using these techniques, the surrogate model could be retrained for patient-specific geometry extracted from MRI. Another key challenge is loading estimate, which is unknown in clinical practice. Dual fluoroscopy combined with multi-body simulations could be used to better assess in vivo loading conditions [69, 70, 23, 71]. Future studies could incorporate a complete functional spinal unit in the FE model, enabling the surrogate to capture IVD mechanics and interactions with surrounding ligaments and facet joints. While the NN offers precise predictions, it lacks uncertainty estimation, which could be useful for assessing unknown loading. Applying stochastic methods in future work could help address this limitation [12]. Incorporating these factors into clinical settings will require effective calibration. Our method facilitates efficient and accurate identification of feasible parameters, representing a key step toward patient-specific simulations.
@@ Acknowledgments
Funding: This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program. Grant no: 101045128-iBack-epic-ERC-2021-COG
@@ B.2 Calibration baselines
The GA [10, 20] was configured with a population size of 20, evolving from initially sampled random configurations within the parameter ranges. The six best-performing individuals from each generation are selected for Crossover and Mutation. During Crossover, four new individuals are created by combining pairs of the selected individuals. Mutation generates four new individuals by randomly altering a parameter of one of the selected individuals. Additionally, Immigration is performed, introducing six new individuals per generation. Every tenth generation, a Restart is performed and 19 new individuals are added to the best individual from the previous generation. The process iterates until an $\overline{\mathcal{R}^{2}}=1$ stopping criteria is met or $120$ generations are reached.
@@ B.3 Models hyperparameters
PR (degree=2)
ridge regularization, alpha=1.0
PR (degree=3)
ridge regularization, alpha=1.0
PR (degree=4)
ridge regularization, alpha=0.11694087489403014
PR (degree=5)
ridge regularization, alpha=0.7741762364391351
@@ Appendix C Detailed results
Table 11: Comparison of interpolation and extrapolation ability of the top performing surrogate models. The results are the mean and standard deviation across the folds.
@@
Sample $C_{10n}$ $C_{01n}$ $C_{10a}$ $k_{1}$ $k_{2}$ $\kappa$ $k_{1c}$ $k_{2c}$ $k_{1r}$ $k_{2r}$ $\alpha$ $\alpha_{c}$ $\alpha_{r}$
Nicolini et al. [39] 1 0.043 0.007 0.060 50.349 199.739 0.268 -0.200 -0.166 -0.041 -0.019 24.054 0.000 0.041
2 0.039 0.048 0.072 38.859 190.633 0.003 -0.054 -0.133 -0.091 -0.026 27.157 0.255 0.156
3 0.117 0.019 0.066 1.556 172.572 0.116 -0.107 -0.025 -0.133 -0.122 26.626 0.015 0.116
4 0.085 0.037 0.067 93.089 181.426 0.267 -0.130 -0.201 -0.187 -0.021 22.690 0.064 0.181
5 0.211 0.009 0.216 104.187 97.086 0.233 -0.166 -0.001 -0.189 -0.159 26.384 0.231 0.023
Heuer et al. [15] Mean 0.062 0.028 0.086 11.800 66.969 0.162 -0.197 -0.006 -0.023 -0.186 38.925 0.017 0.009
Table 13: Result calibrated configurations for experimental data with our method.
Appendix D Running time analysis
PGD w/ NN requires an initial dataset generation step, where 1024 samples are created, each with four FE simulations (one per load case), resulting in 4096 total simulations. Each simulation takes approximately 208 seconds, leading to a total dataset generation time of 236.9 hours or 9.87 days. The model training time is negligible at 1.03 minutes, and once trained, calibration takes only 2.58 seconds, irrespective for the number of specimens.
For a traditional GA setup, using the FE simulations and not the surrogate predictions, we derive worst-case and average-case estimates following the convergence plot in Fig. 13. We assume that 16 (average-case) or 47 (worst-case) generations would be required for the algorithm to converge to an $\overline{\mathcal{R}^{2}}$ higher than 0.9. These result in $20+(1\times 19)+(14\times 14)=235$ or $20+(4\times 19)+(42\times 14)=684$ individuals, with each individual requiring four simulations as before. Here simulations are longer than for the NN, taking 250 seconds (as some specimens involve higher moments), resulting in 2.7 (average-case) or 7.9 (worst-case) days per specimen.
Figure 13: GA w/NN evolution of $\overline{\mathcal{R}^{2}}$ values for experimental specimens. 3rd sample by Nicolini et al. [39] requires more than double the amount of generations to converge to a score higher than 0.9.
Appendix E List of abbreviations